OpenStudy (anonymous):

3y-2 2y-1 _____ + ______ 4 8

OpenStudy (kendricklamar2014):

Still need help?

OpenStudy (anonymous):

yeah

OpenStudy (kendricklamar2014):

Ok lets see....

OpenStudy (candyme):

\[\frac{ 3y-2 }{ 4 }-\frac{ 2y-1 }{ 8 }\] So you just want to combine these, correct?

OpenStudy (anonymous):

yup

OpenStudy (kendricklamar2014):

Step 1. Distribute = \[\frac{ 3 }{ 4 } y + \frac{ -1 }{ 2 } + \frac{ 1 }{ 4 } y + \frac{ -1 }{ 8 }\]

OpenStudy (kendricklamar2014):

Step 2. Combine Like Terms: = \[( \frac{ 3 }{ 4 }y + \frac{ 1 }{ 4 } y) + (\frac{ -1 }{ 2 } + \frac{ -1 }{ 8 })\]

OpenStudy (candyme):

I will let @KendrickLamar2014 finish, then show you a way that may/may not make sense to you

OpenStudy (anonymous):

ok @candyme

OpenStudy (kendricklamar2014):

Which equals: \[y + \frac{ -5 }{ 8 }\] and thats your answer

OpenStudy (anonymous):

@KendrickLamar2014 the answer key shows that it equals  8y-5 _____ 8

OpenStudy (candyme):

So your book is doing it the way I did.\[2(\frac{ 3y-2 }{ 4 })+\frac{ 2y-1 }{ 8 }\] Distribute that 2\[\frac{ 6y-4 }{ 8 } +\frac{ 2y-1 }{ 8 }\] \[\frac{ 6y+2y-4-1 }{ 8 }\]\[\frac{ 8y-5 }{ 8 }\]

OpenStudy (candyme):

@KendrickLamar2014 took it a step further and simplified the 8y/8 part

OpenStudy (candyme):

I multiplied it by 2 in the beginning so I could place all of it on one fraction. I needed a common denominator

OpenStudy (kendricklamar2014):

haha sorry about that @frebrezze

OpenStudy (candyme):

Technically when you get into higher maths you will have to do it his way, but unless it says to simplify all the way I usually leave it